# footnotetext: $\dagger$ First authors

### v1

@@ 7 Discussion
We derive exact solutions to the learning dynamics within a tractable model class: deep linear networks. While our findings extend the range of analytically describable two-layer linear network problems, they are still limited by a set of assumptions. In particular, relaxing the assumptions that input covariance must be white and that initialization must be $\lambda$-balanced could bring the analysis closer to practical applications in machine learning and neuroscience. Moving towards the nonlinear setting would also make the findings more applicable in real-world scenarios. Additionally, removing the requirement for networks to be full rank would enable us to describe the dynamics of linear autoencoders, including settings important in practice, such as LoRA Hu et al. (2022). Despite these limitations, our solutions provide valuable insights into network behaviour. We examine the transition between the rich and lazy regimes by analyzing the dynamics as a function of $\lambda$—the relative scale—across its full range from positive to negative infinity. Our analysis demonstrates that the relative scale, $\lambda$, is pivotal in managing the transition between rich and lazy regimes. Notably, we identify a structured lazy regime that promotes transfer learning. Building on previous work Kunin et al. (2024) that shows these findings extend to basic nonlinear settings and practical scenarios, our theory suggests that further exploration of unbalanced initialization could optimize efficient feature learning. We leave for future work, the extension of this initialization to deep networks. Future work will focus on extending the application of the solution to the dynamics of fine-tuning and linear autoencoders.
@@ Acknowledgments
Author Contributions
Clementine and Nicolas led the work on the equal input-output solution to the exact dynamics, with support from Daniel. Clementine extended this work to the unequal input-output case, examining the impact of the architecture. Clementine and Nicolas also spearheaded the development of the minimal model for rich and lazy learning, emphasizing exact solutions and representation convergence. Daniel, Nicolas, and Clementine collaborated on analyzing the dynamics of singular values, while Lukas led the investigation into representation robustness and sensitivity to noise. Clementine and Nicolas primarily conducted the analysis of the continual learning framework, while Alexandra and Clementine focused on transfer learning in the rich and lazy settings. Clementine took the lead on the reversal learning analysis. She was primarily responsible for leading the manuscript writing, with assistance from Daniel. All authors contributed to writing the appendix and refining the final manuscript.
@@ A.1 Appendix: Balanced Condition
For $\lambda=0$ we have Zero-Balanced given as A5 (). ${\bf W}_{1}(0){\bf W}_{1}(0)^{T}={\bf W}_{2}(0)^{T}{\bf W}_{2}(0)$.
@@ A.2 Discussion Assumptions
Although the whitened input assumption is quite strong, it is commonly used in analytical work to obtain exact solutions, and much of the existing literature relies on these solutions Fukumizu (1998); Braun et al. (2022); Kunin et al. (2024) . Kunin et al. (2024) goes further by exploring the implicit bias of the trajectory without relying on exact solutions. When $X^{\intercal}X$ is low-rank, they can only predict the trajectories in the limit as $\lambda\to\pm\infty$. If the interpolating manifold is one-dimensional, the solution can be solved exactly in terms of $\lambda$ (black dots).

### v2

@@ 7 Discussion
We derive exact solutions to the learning dynamics of deep linear networks. While our findings extend the range of analytically describable two-layer linear network problems, they are still limited by a set of assumptions. In particular, relaxing the assumptions that input covariance must be white and that initialization must be $\lambda$-balanced could bring the analysis closer to practical applications in machine learning and neuroscience. Moving towards the nonlinear setting would also make the findings more applicable in real-world scenarios. Despite these limitations, our solutions provide valuable insights into network behavior. We examine the transition between the rich and lazy regimes by analyzing the dynamics as a function of $\lambda$—the relative scale—across its full range from positive to negative infinity. Our analysis demonstrates that the relative scale, $\lambda$, is pivotal in managing this transition. Notably, we identify a structured lazy regime that promotes transfer learning. Building on previous work (Kunin et al., 2024) that shows these findings extend to basic nonlinear settings and practical scenarios, our theory suggests that further exploration of unbalanced initialization could optimize efficient feature learning. We leave for future work, the extension of this initialization to deep networks. Future work will focus on extending the application of the solution to the dynamics of fine-tuning and linear autoencoders.
@@ A.1 Appendix: Balanced Condition
For $\lambda=0$ we have Zero-Balanced given as A4 (). ${\bf W}_{1}(0){\bf W}_{1}(0)^{T}={\bf W}_{2}(0)^{T}{\bf W}_{2}(0)$.
@@ A.2 Discussion Assumptions
Although the whitened input assumption is quite strong, it is commonly used in analytical work to obtain exact solutions, and much of the existing literature relies on these solutions Fukumizu (1998); Braun et al. (2022); Kunin et al. (2024). While relaxing this assumption prevents the exact description of network dynamics, Kunin et al. (2024) examine the implicit bias of the training trajectory without relying on whitened inputs. If the interpolating manifold is one-dimensional, the solution can be solved exactly in terms of $\lambda$. Their findings demonstrate a similar quantitative dependence on $\lambda$, governing the implicit bias transition between rich and lazy regimes. Furthermore, recent advancements, such as the ”decorrelated backpropagation” technique introduced by Dalm et al. (2024) which whitens inputs during training, showing that optimizing for whitened inputs can actually be done in practice and improve efficiency in real-world applications. Importantly, This study highlights that in certain real-world scenarios, whitening can provide a more optimal learning condition. This approaches emphasize the potential advantages of input whitening for downstream tasks, reinforcing the validity of our assumption.